A company chief financial officer (CFO) is estimating his company’s cost of debt. The company’s bonds offer a coupon rate of 8% with semiannual payments, has exactly 3 years remaining until maturity, and each $1,000 par bond is currently priced at $960.

The company’s cost of debt is closest to:

A) 4.78%

B) 5.48%

C) 8.00%

D) 9.57%
Answer #1

cost of debt  = Yield To Maturity(YTM) = (interest per period+ ((Redemption price - Current market price) / life remaining to maturity)) / ((.4*Redemption price)+ (.6*Current market price))

= ((1000*8%/2)+((1000-960)/(3*2))/(.4*1000+.6*960)

= (40+(40/6))/(400+576)

= (40+6.66666666667)/976

= 46.66666666667/976

= 4.78%